Find the length of the segment that joins P(-1,3) and Q(-2,6) and then find the midpoint.
Answer by checkley77(12844)

P(-1,3) and Q(-2,6)
x distance=-2+1=1
y distance=6-3=3
1^1+3^2=c^2
1+9=c^2
c^2=10c=sqrt10
c=3.162 is the distance between these 2 points.
The midpoint is:
x=(-2-1)/2=-3/2=-1.5
y=(6+3)/2=4.5
Midpoint=(-1.5,4.5)